# A method for sparse and robust independent component analysis

### v2

@@ 3.3 Computational aspects
The computational burden of SICS comes primarily from two parts: computing the scatter matrices $S_{1}$ and $S_{2}$, and running Algorithm 1 for them. The complexity of the former depends greatly on the particular choices of scatters. For example, the classical FOBI-estimate uses the moment-based matrices $S_{1}=\mathrm{Cov}(x)$ and $S_{2}=S_{\mathrm{FOBI}}$ which are very fast to compute even for high $n$ and $p$, whereas the computation of symmetrized robust scatter matrices scales rather badly with both the sample size and data dimension. The runtime of Algorithm 1 comprises of matrix factorizations of low-rank matrices and the updating of $\beta_{j}$ which is carried out using solvebeta [60], a LASSO-solver based on Cholesky decompositions and efficient tracking of inactive and active variables. To demonstrate the computational burden of SICS, we simulated data from the model described later in Section 4 with the choices $n=200,400,600$ and $p=10,20,\ldots,100$ and measured the average time (over 10 replications) it takes to estimate $k=5$ sparse independent components with the sparsity parameter $r=7$. The timing was measured separately for the scatter computation and for Algorithm 1 and we considered two versions of SICS, non-robust (FOBI) and robust (symmetrized Huber and $t_{1}$-MLE).
@@
Fig. 1: Average computation time of SICS as a function of sample size $n$ and data dimension $p$, divided into the computation of the scatters and Algorithm 1.
@@
The results are depicted in Figure 1 and show that Algorithm 1 is the most stable part of SICS timewise. Its runtime appears unaffected by $n$ and scales linearly in the dimension $p$. Also, as expected, the choice of scatters has no impact on this part of the running time. Moving to scatter computation, the FOBI-estimate is extremely fast to compute for all $n,p$, whereas the robust scatters take significantly longer and the growth in their computation time is superlinear in both $n$ and $p$. Thus, for our proposed robust SICS, the main bottleneck are the scatter matrices themselves and, more precisely, the symmetrization operation which is used to obtain the needed independence property and inflates the effective sample size to $n^{2}$, see Section 2. A possible way to alleviate this cost would be to use in the symmetrization only a subset of the pairwise distances [13], trading estimation accuracy for computational speed. Similarly, to manage large values of $p$, one could run SICS repeatedly with several random small subsets of variables to gain a preliminary idea on which features are the most important, pruning the rest, and then running SICS on the resulting data with smaller $p$ to obtain the final estimate. Since the data we study in this work are still manageable in size, we did not resort to these shortcuts.
@@ 4.1 Simulation study #1: Benefits of robustness and sparsity
We investigated the performance of the proposed method SICS with simulations. The task was to estimate the sparse unmixing matrix $\Omega^{-1}$ for a data matrix $X=(x_{1},\ldots,x_{n})^{\prime}$ generated from the IC model $X=Z\Omega^{\prime}$ where $Z=(z_{1},\ldots,z_{n})^{\prime}$. In every iteration, we first generated a random unmixing matrix $\Omega^{-1}\in\mathbb{R}^{p{\times}p}$, where every row has a preset number $q$ of non-zero elements coming from the Uniform([3,1][1,3])([-3,-$\cup[1$,3])-distribution. Then, we generated a random source matrix $Z\in\mathbb{R}^{n{\times}p}$ with the elements of the first column coming from the standardized Laplace distribution, the second column coming from standardized uniform distribution and rest coming from the standard normal distribution. Last, we generated the data matrix $X=Z\Omega^{\prime}$ using $Z$ and $\Omega$. We contaminated the data matrix $X$ by replacing $5\%$ of the rows with observations coming from $N(0,3)$. All simulations were also run similarly to the non-contaminated datasets to evaluate the effect of the contamination to the estimators. We used different variations of proposed SICS algorithm to estimate the coefficient vector $\beta_{1}$ of the first independent component and calculated its absolute distance to the true first row of the matrix $\Omega$. Because the sign of the coefficient vector is arbitrary, distance to vector with flipped signs was used if it was smaller. Then we took the median of this over $N$ iterations.
The methods to compare were SICS with four different numbers of estimated non-zero coefficients $r$: full $p$ (non-sparse estimate), true $q$ (oracle estimate), $q+3$ and “forced q+3q+3” where all $p$ coefficients were first estimated in a non-sparse way and then the pq3p-q-3 with the smallest absolute values were set to zero (making it a thresholding estimate). All of those were run with a non-robust pair of scatter matrices, the covariance matrix and the FOBI matrix, and a robust pair, symmetrized $t$-distribution based ($\nu=1$) and symmetrized Huber’s $M$-estimator (where $c$ is chosen so that $\mathbb{P}(\chi^{2}_{p}\leq c^{2}/2)=0.9$).
@@ 4.3 Simulation study #3: Behavior under increasing $p$
The results are shown as a function of $p$ in Figure 5. We first note that the scenario is quite challenging as the number of data points is kept fixed while at the same time we introduce more and more noise to the model. Thus, it is expected that the error rate is increasing in $p$. However, we observe that, despite the wrong choice of $r$, this increase is for SICS much milder than for ICS, showing that the larger the dimension, the more we benefit from sparse estimation. A prospective future task is to study the theoretical growth rate of the error in scenarios where $n,p_{n}\rightarrow\infty$. For example, for regular LASSO one can achieve, under specific conditions, convergence rate of the order $\{(\log p)/n\}^{1/2}$, see [22, Section 11]. Finally, interpreting the results from an efficiency viewpoint, we observe from Figure 5, e.g., that SICS needs only $n=1000$ observations to achieve equal or better performance than ICS with $n=2000$.
@@ 4.4 Simulation study #4: Comparison with competing methods
Next, we compare SICS with some other approaches to doing sparse ICA. A simple way of obtaining a sparse version from the implementations of the well-known ICA methods, JADE and FastICA (in R-packages JADE [35] and fICA [36]), is to use thresholding to put a subset of the coefficients to zero. Let $r$ be the parameter giving the estimated number of non-zero coefficients. In hard thresholding prp-r smallest coefficients are set to zero after estimation. In soft thresholding, in addition to that, all other coefficients are shrunk by the absolute value of the largest coefficient among those that were set to zero. In addition to these four competitors, we also include a sparse version of FOBI achieved through taking the sparse generalized eigendecomposition of covariance matrix and FOBI-matrix using the implementation of RIFLE [51] in [50], with the tuning parameter values $K=1$ and $\lambda=2\sqrt{\log p/n}$ in initial.convex.
@@
Fig. 6: The results of a comparison between SICS and competitors
@@ 5 Robust causal discovery
The obtained causal graphs are shown in Figure 7 and show that both methods more or less agree on the causal structure of the data, indirectly indicating that the data are not likely to contain significant outliers (which likely stems from the fact that the data represent a curated clinical study). The main feature of interest is that most of the 10 explanatory variables are estimated to be causes of the disease progression variable DP. Additionally, a relation between the covariates LDL and TC was discovered. We note that these plots should not be interpreted as the full causal graphs between the 11 variables, but rather as estimates of the set of strongest causal relationships between the variables.
We next take these findings as the ground truth, and continue the experiment by creating a contaminated version of the data set, obtained by randomly selecting $10\%$ of the subjects and replacing all their measurements with i.i.d. Gaussian noise with standard deviation $\sigma=20$. We then applied the same causal discovery approach to this contaminated data set, with the hopes of still being able to find the relevant structure, despite the contamination. The bootstrapped causal graphs estimated from the contaminated data are shown in Figure 8 and reveal that the structure of the non-robust graph has completely changed, with only two of the original edges remaining. That is, the outliers have made it impossible for standard, non-robust ICA to find almost any causal relations. However, the graph from robust and sparse ICA retains most of the connections, in particular the edges between the disease progression and the covariates AGE, SEX, BMI, MAP, TC, LDL and GLU. We thus conclude that robust SICS offers a reliable and outlier-resistant method for linear causal discovery.
@@ Acknowledgments
The authors are grateful to the anonymous reviewer whose comments were of great help in improving the manuscript. The authors would like to thank Henri Nyberg for his constructive comments and feedback on the manuscript. The work of LH and JV was supported by the Research Council of Finland (grants 347501, 353769). The authors would like to thank Andreas Artemiou who brought the work by [31] to their attention.
@@ Appendix B Proof of Theorem 4
We assume that $M_{n},G_{n}$ are symmetric $p\times p$ matrices such that
$M_{n}\rightarrow_{p}M,\quad G_{n}\rightarrow_{p}G,$
as $n\rightarrow\infty$, where $M,G$ are symmetric and positive definite $p\times p$ matrices. The first, second and last eigenvalues of $G_{n}^{-1/2}M_{n}G_{n}^{-1/2}$ are denoted in the following by $\rho_{n},\psi_{n},\Omega_{n}$, respectively. The population counterparts of these are denoted by $\rho,\psi,\Omega$ and we assume that $\rho>\psi$ and $\Omega>0$. The population and sample eigenvectors corresponding to the leading eigenvalues $\rho_{n},\rho$ are denoted by $u_{n},u\in\mathbb{R}^{p}$, respectively. Finally, we denote the smallest eigenvalues of $G_{n}$ and $G$ as $\alpha_{n}$ and $\alpha>0$, respectively.
@@
For $a,b\in\mathbb{R}^{p}$, $\|a\|_{2}=1$, we have
$\sum_{i=1}^{p}\|G^{-1/2}r_{i}-ab^{\prime}r_{i}\|_{2}^{2}=\mathrm{tr}(G^{-1/2}MG^{-1/2})-2b^{\prime}MG^{-1/2}a+b^{\prime}Mb,$
where $(r_{1},\ldots,r_{p})=M^{1/2}$.
@@
The minimizer $(\alpha,\beta)$ of
$(a,b)\mapsto\sum_{i=1}^{p}\|G^{-1/2}r_{i}-ab^{\prime}r_{i}\|_{2}^{2},$
over $a,b\in\mathbb{R}^{p}$, $\|a\|_{2}=1$ is
$\alpha=u\quad\mbox{and}\quad\beta=G^{-1/2}u.$
@@
    Version I:
    $(a_{n,\lambda_{n}},b_{n,\lambda_{n}})=\mathrm{argmin}f_{n,\lambda_{n}}(a,b),\quad a,b\in\mathbb{R}^{p},\|a\|_{2}=1,$
    where $f_{n,\lambda_{n}}(a,b)=-2b^{\prime}M_{n}G_{n}^{-1/2}a+b^{\prime}M_{n}b+\lambda_{n}\|b\|_{1}$.
@@
    Version II:
    $(a_{n,0},b_{n,0})=\mathrm{argmin}f_{n,0}(a,b),\quad a,b\in\mathbb{R}^{p},\|a\|_{2}=1,$
    where $f_{n,0}$ is as in Version I above.
@@
    Version III:
    $(\alpha,\beta)=\mathrm{argmin}f(a,b),\quad a,b\in\mathbb{R}^{p},\|a\|_{2}=1,$
    where $f(a,b)=-2b^{\prime}MG^{-1/2}a+b^{\prime}Mb$.
@@
For a vector $m\in\mathbb{R}^{p}$, we define $a^{*}_{n,\lambda_{n},m}$ to be the minimizer of the objective function
$a\mapsto f_{n,\lambda_{n}}(a,b_{n,0}+m),$
over $a\in\mathbb{R}^{p},\|a\|_{2}=1$. As the objective function is symmetric in the sense that $f_{n,\lambda_{n}}(a,b)=f_{n,\lambda_{n}}(-a,-b)$, it is sufficient to restrict our attention to vectors $m\in\mathbb{R}^{p}$ such that $b_{n,0}^{\prime}G_{n}(b_{n,0}+m)\geq 0$ (corresponding to a certain half-space of $\mathbb{R}^{p}$).
@@
For any $m\in\mathbb{R}^{p}$ such that $b_{n,0}^{\prime}G_{n}(b_{n,0}+m)\geq 0$, we have
$f_{n,\lambda_{n}}(a^{*}_{n,\lambda_{n},m},b_{n,0}+m)-f_{n,\lambda_{n}}(a^{*}_{n,\lambda_{n},0},b_{n,0})$
$\geq$ $\min\{C_{1n},C_{2n}\}\|A_{n}^{-1}\|_{2}^{-2}\|G_{n}^{-1/2}\|_{2}^{-2}\|m\|_{2}^{2}-\lambda_{n}\sqrt{p}\|m\|_{2},$
where
$C_{1n}$ $:=0.4\rho_{n}\{\rho_{2}(A_{n})^{-1}-1\},$
$C_{2n}$ $:=(\sqrt{0.6}-\sqrt{0.4})^{2}\rho_{n},$
and $A_{n}:=G_{n}^{-1/2}M_{n}G_{n}^{-1/2}/\rho_{n}$.
@@
$\begin{split}&f_{n,\lambda_{n}}(a^{*}_{n,\lambda_{n},m},b_{n,0}+m)-f_{n,\lambda_{n}}(a^{*}_{n,\lambda_{n},0},b_{n,0})\\ =&-2\|\rho_{n}u_{n}+M_{n,0}G_{n}^{1/2}m\|_{2}+2\rho_{n}+2b_{n,0}^{\prime}M_{n}m+m^{\prime}M_{n}m\\ &+\lambda_{n}\{\|b_{n,0}+m\|_{1}-\|b_{n,0}\|_{1}\},\end{split}$ (3)
where $M_{n,0}:=G_{n}^{-1/2}M_{n}G_{n}^{-1/2}$ and $\rho_{n}$ denotes the largest eigenvalue of this matrix. Let now $A_{n}:=M_{n,0}/\rho_{n}$ and express $m$ as $m=G_{n}^{-1/2}A_{n}^{-1}v$ for some $v\in\mathbb{R}^{p}$. Our earlier assumption that $b_{n,0}^{\prime}G_{n}(b_{n,0}+m)\geq 0$ then takes the form
$0\leq$ $b_{n,0}^{\prime}G_{n}(b_{n,0}+m)$
$=$ $u_{n}^{\prime}(u_{n}+G_{n}^{1/2}m)$
$=$ $1+u_{n}^{\prime}A_{n}^{-1}v$
$=$ $1+u_{n}^{\prime}v,$
@@
$-2\|\rho_{n}u_{n}+M_{n,0}G_{n}^{1/2}m\|_{2}+2\rho_{n}+2b_{n,0}^{\prime}M_{n}m+m^{\prime}M_{n}m$
$=$ $\rho_{n}\{-2\|u_{n}+A_{n}G_{n}^{1/2}m\|_{2}+2+2u_{n}^{\prime}A_{n}G_{n}^{1/2}m+m^{\prime}G_{n}^{1/2}A_{n}G_{n}^{1/2}m\}$
$=$ $\rho_{n}\{-2\|u_{n}+v\|_{2}+2+2u_{n}^{\prime}v+v^{\prime}A_{n}^{-1}v\}.$
Then, have the following identity
$(\|u_{n}+v\|_{2}-1)^{2}=$ $1+2u_{n}^{\prime}v+v^{\prime}v-2\|u_{n}+v\|_{2}+1.$
Hence, we can write
$\begin{split}&-2\|\rho_{n}u_{n}+M_{n,0}G_{n}^{1/2}m\|_{2}+2\rho_{n}+2b_{n,0}^{\prime}M_{n}m+m^{\prime}M_{n}m\\ =&\rho_{n}\{(\|u_{n}+v\|_{2}-1)^{2}+v^{\prime}(A_{n}^{-1}-I_{p})v\}.\end{split}$ (4)
Now, as $A_{n}$ has its eigenvalues in $[0,1]$, then $A_{n}^{-1}-I_{p}$ has eigenvalues in $0,\infty)$ with the eigenvector $u_{n}$ corresponding to the eigenvalue 0. Decompose then $v$ such that $v=P_{n}v+Q_{n}v$ where $P_{n}:=u_{n}u_{n}^{\prime}$ is the projection onto $u_{n}$ and $Q_{n}:=I_{p}-P_{n}$. As such, $P_{n}(A_{n}^{-1}-I_{p})=0$ and we have
$v^{\prime}(A_{n}^{-1}-I_{p})v=$ $v^{\prime}Q_{n}(A_{n}^{-1}-I_{p})Q_{n}v$
$=$ $\|Q_{n}v\|_{2}^{2}\frac{(Q_{n}v)^{\prime}}{\|Q_{n}v\|_{2}}(A_{n}^{-1}-I_{p})\frac{Q_{n}v}{\|Q_{n}v\|_{2}}$
$\geq$ $\|Q_{n}v\|_{2}^{2}\rho_{p-1}(A_{n}^{-1}-I_{p})$
$=$ $\|Q_{n}v\|_{2}^{2}\{\rho_{2}(A_{n})^{-1}-1\},$

### v3

@@ 3.4 Computational aspects
The computational burden of SICS comes from computing the scatter matrices $S_{1}$ and $S_{2}$ and running Algorithm 1 for them, the latter comprising of multiple steps. The complexity of the scatter computation depends greatly on the particular choices of scatters. For example, the classical FOBI-estimate uses the moment-based matrices $S_{1}=\mathrm{Cov}(x)$ and $S_{2}=S_{\mathrm{FOBI}}$ which are very fast to compute even for high $n$ and $p$, whereas the computation of symmetrized robust scatter matrices scales rather badly with both the sample size and data dimension. The runtime of Algorithm 1 is divided to (a) computing the standard ICS-solution, (b) updating the $\beta_{j}$ using solvebeta [61], a LASSO-solver based on Cholesky decompositions and efficient tracking of inactive and active variables, and (c) matrix factorizations of low-rank matrices. To demonstrate the computational burden of SICS, we simulated data from the model described later in Section 4 with the choices $n=200,400,600$ and $p=10,20,\ldots,100$ and measured the average time (over 20 replications) it takes to estimate $k=5$ sparse independent components with the sparsity parameter $r=7$. The full timing was divided into four parts: scatter computation, lines 1–3 of Algorithm 1 (standard ICS), lines 5–9 (LASSO-update) and lines 10–13 (decompositions). We considered two versions of SICS, non-robust (FOBI) and robust (symmetrized Huber and $t_{1}$-MLE). The timings were measured on a 14” MacBook Pro M3 with 16 GB memory.
@@
Fig. 1: Average computation time of SICS as a function of sample size $n$ and data dimension $p$, divided into the computation of the scatters and the different parts of Algorithm 1. The $y$-axis has a logarithmic scale.
The results are depicted in Figure 1 and show that the main bottleneck of SICS is the computation of robust scatters which scales superlinearly in both $n$ and $p$ (red dashed line). This is caused by the symmetrization operation which is needed for the independence property and inflates the effective sample size to $n^{2}$, see Section 2. In comparison, computing the FOBI-estimate is extremely fast for all $n,p$. As Algorithm 1 consists of optimization routines and matrix decompositions for the pre-computed $p\times p$ scatters $S_{1},S_{2}$, it is natural that its computation time depends very little on $n$ and the choice of scatters. Of the three components of Algorithm 1, the computation of the standard ICS-solution (lines 1-3) is the fastest part. This is as expectable since the other two components require iteration (the repeat-until loop). Computing the matrix decomposition (lines 10-13) is an order of magnitude faster than the LASSO-update (5-9). This is because the matrices in question are low-rank, having sizes $p\times k$ and $k\times k$ with $k=5$ here, leading to efficient computation of the decompositions.
@@
A possible way to alleviate the computation cost of the robust scatters would be to use in the symmetrization only a subset of the pairwise distances [13], trading estimation accuracy for computational speed. Similarly, to manage large values of $p$, one could run SICS repeatedly with several random small subsets of variables to gain a preliminary idea on which features are the most important, pruning the rest, and then running SICS on the resulting data with smaller $p$ to obtain the final estimate. Since the data we study in this work are still manageable in size, we did not resort to these shortcuts.
@@ 4.1 Simulation study #1: Benefits of robustness and sparsity
We investigated the performance of the proposed method SICS with simulations. The task was to estimate the sparse unmixing matrix $\Omega^{-1}$ for a data matrix $X=(x_{1},\ldots,x_{n})^{\top}$ generated from the IC model $X=Z\Omega^{\top}$ where $Z=(z_{1},\ldots,z_{n})^{\top}$. In every iteration, we first generated a random unmixing matrix $\Omega^{-1}\in\mathbb{R}^{p{\times}p}$, where every row has a preset number $q$ of non-zero elements coming from the Uniform([3,1][1,3])([-3,-$\cup[1$,3])-distribution. Then, we generated a random source matrix $Z\in\mathbb{R}^{n{\times}p}$ with the elements of the first column coming from the standardized Laplace distribution, the second column coming from standardized uniform distribution and rest coming from the standard normal distribution. Last, we generated the data matrix $X=Z\Omega^{\top}$ using $Z$ and $\Omega$. We contaminated the data matrix $X$ by replacing $5\%$ of the rows with observations coming from $N(0,3)$. All simulations were also run similarly to the non-contaminated datasets to evaluate the effect of the contamination to the estimators. We used different variations of proposed SICS algorithm to estimate the coefficient vector $\beta_{1}$ of the first independent component and calculated its absolute distance to the true first row of the matrix $\Omega$. Because the sign of the coefficient vector is arbitrary, distance to vector with flipped signs was used if it was smaller. Then we took the median of this over $N$ iterations.
The methods to compare were SICS with four different numbers of estimated non-zero coefficients $r$: full $p$ (non-sparse estimate), true $q$ (oracle estimate), $q+3$ and “hard q+3q+3” where all $p$ coefficients were first estimated in a non-sparse way and then the pq3p-q-3 with the smallest absolute values were set to zero (making it a hard thresholding estimate). All of those were run with a non-robust pair of scatter matrices, the covariance matrix and the FOBI matrix, and a robust pair, symmetrized $t$-distribution based ($\nu=1$) and symmetrized Huber’s $M$-estimator (where $c$ is chosen so that $\mathbb{P}(\chi^{2}_{p}\leq c^{2}/2)=0.9$).
@@ 4.3 Simulation study #3: Behavior under increasing $p$
The results are shown as a function of $p$ in Figure 5. We first note that the scenario is quite challenging as the number of data points is kept fixed while at the same time we introduce more and more noise to the model. Thus, it is expected that the error rate is increasing in $p$. However, we observe that, despite the wrong choice of $r$, this increase is for SICS much milder than for ICS, showing that the larger the dimension, the more we benefit from sparse estimation. A prospective future task is to study the theoretical growth rate of the error in scenarios where $n,p_{n}\rightarrow\infty$. For example, for regular LASSO one can achieve, under specific conditions, convergence rate of the order $\{(\ln p)/n\}^{1/2}$, see [22, Section 11]. Finally, interpreting the results from an efficiency viewpoint, we observe from Figure 5, e.g., that SICS needs only $n=1000$ observations to achieve equal or better performance than ICS with $n=2000$.
@@ 4.4 Simulation study #4: Comparison with competing methods
Next, we compare SICS with some other approaches to doing sparse ICA. A simple way of obtaining a sparse version from the implementations of the well-known ICA methods, JADE and FastICA (in R-packages JADE [35] and fICA [36]), is to use thresholding to put a subset of the coefficients to zero. Let $r$ be the parameter giving the estimated number of non-zero coefficients. In hard thresholding prp-r smallest coefficients are set to zero after estimation. In soft thresholding, in addition to that, all other coefficients are shrunk by the absolute value of the largest coefficient among those set to zero. In addition to these four competitors, we also include a sparse version of FOBI achieved through taking the sparse generalized eigendecomposition of covariance matrix and FOBI-matrix using the implementation of RIFLE [52] in [51], with the tuning parameter values $K=1$ and $\lambda=2\sqrt{\ln p/n}$ in initial.convex.
@@
The median computation times (over 80 repetitions) are presented in Table 1. The computations were run on Dell Latitude 5520 15,6” i5 with 8GB memory. The timings correspond to case $r=q$, but the results are similar for other values of $r$. It seems that for all method except fICA, the computation times are not dependent on the sample size. This is because the other methods operate directly on $p\times p$ matrices that are computed only once, whereas fICA requires repeated evaluations of objective function gradients separately for each observation.
$n$ 500 1000 1500 2000 2500
SICS 0.448 (0.152) 0.495 (0.123) 0.501 (0.144) 0.488 (0.198) 0.505 (0.229)
fICA 0.780 (0.047) 1.309 (0.095) 1.952 (0.267) 2.440 (0.187) 3.042 (0.312)
JADE 0.045 (0.012) 0.072 (0.012) 0.092 (0.011) 0.117 (0.013) 0.143 (0.019)
RIFLE 0.089 (0.116) 0.059 (0.078) 0.069 (0.092) 0.072 (0.094) 0.063 (0.083)
Table 1: Median running times in seconds (with MAD in parenthesis) for different methods and sample sizes in simulation #4.
@@ 5 Robust causal discovery
The obtained causal graphs are shown in Figure 7 and show that both methods more or less agree on the causal structure of the data, indirectly indicating that the data are not likely to contain significant outliers (which likely stems from the fact that the data represent a curated clinical study). The main feature of interest is that most of the 10 explanatory variables are estimated to be causes of the disease progression variable DP. Additionally, a relation between the covariates LDL and TC was discovered. We note that these plots should not be interpreted as the full causal graphs between the 11 variables, but rather as estimates of the set of strongest causal relationships between the variables. We also computed the runtimes (14” MacBook Pro M3, 16 GB memory) of producing the causal graph both for the non-robust and robust method and these were 11 s and 6 min 44 s, respectively. The difference in the orders of magnitude is in line with the simulation results in Section 3.4.
We next take these findings as the ground truth, and continue the experiment by creating a contaminated version of the data set, obtained by randomly selecting $10\%$ of the subjects and replacing all their measurements with i.i.d. Gaussian noise with standard deviation $\sigma=20$. We then applied the same causal discovery approach to this contaminated data set, with the hopes of still being able to find the relevant structure, despite the contamination. The bootstrapped causal graphs estimated from the contaminated data are shown in Figure 8 and reveal that the structure of the non-robust graph has completely changed, with only two of the original edges remaining. That is, the outliers have made it impossible for standard, non-robust ICA to find almost any causal relations. However, the graph from robust and sparse ICA retains most of the connections, in particular the edges between the disease progression and the covariates AGE, SEX, BMI, MAP, TC, LDL and GLU. We thus conclude that robust SICS offers a reliable and outlier-resistant method for linear causal discovery. The runtimes for the contaminated data were 10 s and 12 min 13 s, for the non-robust and robust approach, respectively. The time difference in the robust method for these and the non-contaminated data (12 min 13 s vs. 6 min 44 s) comes from the computation of the symmetrized Huber’s M-estimator: the number of pairs with long distances between them is large in the contaminated data, and when these get truncated to constant size by Huber’s weight function [50], one obtains a less steep gradient and a larger number of iterations is needed for convergence.
@@ Acknowledgments
The authors are grateful to the anonymous reviewers whose comments were of great help in improving the manuscript. The authors would like to thank Henri Nyberg for his constructive comments and feedback on the manuscript. The work of LH and JV was supported by the Research Council of Finland (grants 347501, 353769, 368494). The authors would like to thank Andreas Artemiou who brought the work by [31] to their attention.
@@ Appendix B Proof of Theorem 4
We assume that $M_{n},G_{n}$ are symmetric $p\times p$ matrices such that $M_{n}\rightarrow_{p}M,\quad G_{n}\rightarrow_{p}G$, as $n\rightarrow\infty$, where $M,G$ are symmetric and positive definite $p\times p$ matrices. The first, second and last eigenvalues of $G_{n}^{-1/2}M_{n}G_{n}^{-1/2}$ are denoted in the following by $\rho_{n},\psi_{n},\Omega_{n}$, respectively. The population counterparts of these are denoted by $\rho,\psi,\Omega$ and we assume that $\rho>\psi$ and $\Omega>0$. The population and sample eigenvectors corresponding to the leading eigenvalues $\rho_{n},\rho$ are denoted by $u_{n},u\in\mathbb{R}^{p}$, respectively. Finally, we denote the smallest eigenvalues of $G_{n}$ and $G$ as $\alpha_{n}$ and $\alpha>0$, respectively.
@@
For $a,b\in\mathbb{R}^{p}$, $\|a\|=1$, we have $\sum_{i=1}^{p}\|G^{-1/2}r_{i}-ab^{\top}r_{i}\|^{2}=\mathrm{tr}(G^{-1/2}MG^{-1/2})-2b^{\top}MG^{-1/2}a+b^{\top}Mb$, where $(r_{1},\ldots,r_{p})=M^{1/2}$.
@@
The minimizer $(\alpha,\beta)$ of $(a,b)\mapsto\sum_{i=1}^{p}\|G^{-1/2}r_{i}-ab^{\top}r_{i}\|^{2}$ over $a,b\in\mathbb{R}^{p}$, $\|a\|=1$ is $\alpha=u$ and $\beta=G^{-1/2}u$.
@@
    Version I: $(a_{n,\lambda_{n}},b_{n,\lambda_{n}})=\mathrm{argmin}f_{n,\lambda_{n}}(a,b)$, $a,b\in\mathbb{R}^{p},\|a\|=1$, where $f_{n,\lambda_{n}}(a,b)=-2b^{\top}M_{n}G_{n}^{-1/2}a+b^{\top}M_{n}b+\lambda_{n}\|b\|_{1}$.
@@
    Version II: $(a_{n,0},b_{n,0})=\mathrm{argmin}f_{n,0}(a,b)$, $a,b\in\mathbb{R}^{p},\|a\|=1$, where $f_{n,0}$ is as in Version I above.
@@
    Version III: $(\alpha,\beta)=\mathrm{argmin}f(a,b)$, $a,b\in\mathbb{R}^{p},\|a\|=1$, where $f(a,b)=-2b^{\top}MG^{-1/2}a+b^{\top}Mb$.
@@
For a vector $m\in\mathbb{R}^{p}$, we define $a^{*}_{n,\lambda_{n},m}$ to be the minimizer of the objective function $a\mapsto f_{n,\lambda_{n}}(a,b_{n,0}+m)$ over $a\in\mathbb{R}^{p},\|a\|=1$. As the objective function is symmetric in the sense that $f_{n,\lambda_{n}}(a,b)=f_{n,\lambda_{n}}(-a,-b)$, it is sufficient to restrict our attention to vectors $m\in\mathbb{R}^{p}$ such that $b_{n,0}^{\top}G_{n}(b_{n,0}+m)\geq 0$ (corresponding to a certain half-space of $\mathbb{R}^{p}$).
@@
For any $m\in\mathbb{R}^{p}$ such that $b_{n,0}^{\top}G_{n}(b_{n,0}+m)\geq 0$, we have
$f_{n,\lambda_{n}}(a^{*}_{n,\lambda_{n},m},b_{n,0}+m)-f_{n,\lambda_{n}}(a^{*}_{n,\lambda_{n},0},b_{n,0})\geq\min\{C_{1n},C_{2n}\}\|A_{n}^{-1}\|_{2}^{-2}\|G_{n}^{-1/2}\|_{2}^{-2}\|m\|^{2}-\lambda_{n}\sqrt{p}\|m\|,$
where $C_{1n}:=0.4\rho_{n}\{\rho_{2}(A_{n})^{-1}-1\}$, $C_{2n}:=(\sqrt{0.6}-\sqrt{0.4})^{2}\rho_{n}$ and $A_{n}:=G_{n}^{-1/2}M_{n}G_{n}^{-1/2}/\rho_{n}$.
@@
$\begin{split}&f_{n,\lambda_{n}}(a^{*}_{n,\lambda_{n},m},b_{n,0}+m)-f_{n,\lambda_{n}}(a^{*}_{n,\lambda_{n},0},b_{n,0})\\ =&-2\|\rho_{n}u_{n}+M_{n,0}G_{n}^{1/2}m\|+2\rho_{n}+2b_{n,0}^{\top}M_{n}m+m^{\top}M_{n}m+\lambda_{n}\{\|b_{n,0}+m\|_{1}-\|b_{n,0}\|_{1}\},\end{split}$ (3)
where $M_{n,0}:=G_{n}^{-1/2}M_{n}G_{n}^{-1/2}$ and $\rho_{n}$ denotes the largest eigenvalue of this matrix. Let now $A_{n}:=M_{n,0}/\rho_{n}$ and express $m$ as $m=G_{n}^{-1/2}A_{n}^{-1}v$ for some $v\in\mathbb{R}^{p}$. Our earlier assumption that $b_{n,0}^{\top}G_{n}(b_{n,0}+m)\geq 0$ then takes the form
$0\leq$ $b_{n,0}^{\top}G_{n}(b_{n,0}+m)=u_{n}^{\top}(u_{n}+G_{n}^{1/2}m)=1+u_{n}^{\top}A_{n}^{-1}v=1+u_{n}^{\top}v,$
@@
$-2\|\rho_{n}u_{n}+M_{n,0}G_{n}^{1/2}m\|+2\rho_{n}+2b_{n,0}^{\top}M_{n}m+m^{\top}M_{n}m=$ $\rho_{n}\{-2\|u_{n}+A_{n}G_{n}^{1/2}m\|+2+2u_{n}^{\top}A_{n}G_{n}^{1/2}m+m^{\top}G_{n}^{1/2}A_{n}G_{n}^{1/2}m\}$
$=$ $\rho_{n}\{-2\|u_{n}+v\|+2+2u_{n}^{\top}v+v^{\top}A_{n}^{-1}v\}.$
Next, we have the identity $(\|u_{n}+v\|-1)^{2}=1+2u_{n}^{\top}v+v^{\top}-2\|u_{n}+v\|+1$. Hence, we can write
$\begin{split}-2\|\rho_{n}u_{n}+M_{n,0}G_{n}^{1/2}m\|+2\rho_{n}+2b_{n,0}^{\top}M_{n}m+m^{\top}M_{n}m=\rho_{n}\{(\|u_{n}+v\|-1)^{2}+v^{\top}(A_{n}^{-1}-I_{p})v\}.\end{split}$ (4)
Now, as $A_{n}$ has its eigenvalues in $[0,1]$, then $A_{n}^{-1}-I_{p}$ has eigenvalues in $0,\infty)$ with the eigenvector $u_{n}$ corresponding to the eigenvalue 0. Decompose then $v$ such that $v=P_{n}v+Q_{n}v$ where $P_{n}:=u_{n}u_{n}^{\top}$ is the projection onto $u_{n}$ and $Q_{n}:=I_{p}-P_{n}$. As such, $P_{n}(A_{n}^{-1}-I_{p})=0$ and we have
$v^{\top}(A_{n}^{-1}-I_{p})v=$ $v^{\top}Q_{n}(A_{n}^{-1}-I_{p})Q_{n}v=\|Q_{n}v\|^{2}\frac{(Q_{n}v)^{\top}}{\|Q_{n}v\|}(A_{n}^{-1}-I_{p})\frac{Q_{n}v}{\|Q_{n}v\|}\geq\|Q_{n}v\|^{2}\rho_{p-1}(A_{n}^{-1}-I_{p})=\|Q_{n}v\|^{2}\{\rho_{2}(A_{n})^{-1}-1\},$